OpenStudy (anonymous):

A triangle has vertices whose coordinates are A (0, 0), B(5, 7), and C(3, 9). Find the coordinates of the midpoints of each side. Midpoint of AB = ( , ) Midpoint of BC = ( , ) Midpoint of CA = ( , )

OpenStudy (anonymous):

AB=(5/2,7/2) BC=(4,8) CA=(3/2,9/2)

OpenStudy (ash2326):

Midpoint of AB= ((0+5)/2, (0+7)/2)=(2.5, 3.5) midpoint of BC= ((5+3)/2, (7+9)/2)=(4, 8) Midpoint of CA= ((0+3)/2,(0+9)/2)=(1.5, 4.5)

OpenStudy (anonymous):

thnx quys